Amanda and Fero are buying a house on a 30-year mortgage. They can only pay $800...

  1. Amanda and Fero are buying a house on a 30-year mortgage. They can only pay $800 per month for a mortgage. If they have an APR of 3.75%, what is the maximum price of a mortgage that they can take out? How much interest will they pay over the 30 year period?
Answer #1

Solution:

Monthly interest rate = 3.75% / 12 = 0.3125%

Nos of monthly periods = 30*12 = 360 periods

Monthly mortgage payment = $800

Maximum price of a mortgage that they can take out = $800 * Cumulative PV factor at 0.3125% for 360 periods

= $800 * 215.92881 = $172,743

Total payment over 30 years = 800*360 = $288,000

Interest to be paid over 30 years = $288,000 - $172,743 = $115.257